### v7

@@ 1 Introduction
Let $\mathcal{O}$ be a bounded convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator with homogeneous Dirichlet boundary conditions. We consider the following stochastic Allen-Cahn equation:
$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
$y(0)=v,$ (1.1b)
@@
where the initial value $v$ is given, and $W_{H}$ is an infinite-dimensional $H$-cylindrical Brownian motion.
@@
The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [4, 10, 15, 27, 33]. They are also an important class of nonlinear stochastic partial differential equations from a mathematical perspective. Recently, researchers in the field of numerical analysis have shown considerable interest in this topic; see [2, 3, 5, 6, 9, 14, 20, 21, 23, 27, 31, 36] and the references therein. We briefly summarize some closely related works of numerical analysis of three-dimensional stochastic Allen-Cahn equations as follows. Feng et al. [9] analyzed finite element approximations of a stochastic Allen-Cahn equation with gradient-type multiplicative noise. Majee and Prohl [23] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative finite-dimensional Wiener process. Bréhier et al. [5] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Jentzen and Pusnik [14] proposed a numerical method for approximating a class of semilinear stochastic evolution equations with non-globally Lipschitz continuous nonlinearities. Qi and Wang [31] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Liu and Qiao [21] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations driven by a multiplicative infinite-dimensional Wiener process. Qi et al. [32] proposed an efficient approximation method to the stochastic Allen-Cahn equation with random diffusion coefficient field and multiplicative noise.
The main motivation of this work includes two components:
  • To the best of our knowledge, existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data. This scenario is of practical significance, as initial phases may include randomly distributed particles. The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis compared to cases with smooth initial data. Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [18, 19]). Secondly, the singularity introduced by non-smooth initial data significantly complicates the process of numerical analysis.
@@
  • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm. However, error estimates using the more general spatial $L^{q}$-norms are relatively scarce.
The main contributions of this work are as follows:
@@
  • For the model problem Eq. 1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are crucial for the numerical analysis. Specifically, we affirm that
    $y\in L^{p}(\Omega;C([0,T];L^{q})),\quad\forall p,q\in(2,\infty).$
@@
    Our analysis uses a suite of primary tools, including: (i) stochastic integration and Itô’s formula in UMD spaces [7]; (ii) the stochastic maximal $L^{p}$-regularity estimates developed in [25, 26]; (iii) the theory of stochastic partial differential equations with Lyapunov conditions from [18, 19]; (iv) the maximal $L^{p}$-regularity theory for deterministic parabolic equations [37].
  • To handle the singularity caused by the non-smooth initial data, by decomposing the error into a singular part and a relatively non-singular part, we derive the error estimate
@@
    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\big{(}h^{\alpha/p}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],$
    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [35, Theoerem 3.2].
  • Given smooth initial data $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we mainly use the standard theory of finite element methods (see, e.g., [1, 35]), the discrete maximal $L^{p}$-regularity [16], and the discrete stochastic maximal $L^{p}$-regularity [17] to derive two rigorous error estimates within the framework of general spatial $L^{q}$-norms. Specifically, for arbitrary parameters $p\in(2,\infty)$ and $q\in[2,\infty)$, and for any chosen $\epsilon>0$, we establish the following error estimates:
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$
@@
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
To achieve a second-order spatial convergence rate, prior numerical analyses in the literature have typically enforced stringent regularity assumptions on the diffusion coefficients. In contrast, this study imposes relatively moderate regularity conditions on the diffusion coefficient $F$ (see Section 2). While our numerical analysis assumes deterministic initial data, extending the results to accommodate random initial data is straightforward.
@@
The remainder of this paper is organized as follows. Section 2 introduces the notational conventions and definitions that are used throughout the paper. Section 3 presents the main results of this work, including the regularity properties of the mild solution to model problem 1.1 with non-smooth initial data, and establishes the convergence of the finite element-based spatial semidiscretization for both non-smooth and smooth initial conditions. A rigorous proof of these results is provided in Section 4. Finally, Section 5 concludes the paper by summarizing the key findings of this study.
@@ 2 Preliminaries
Notational Conventions. Throughout this paper, we will use the following conventions. For Banach spaces $E_{1}$ and $E_{2}$, the space of all bounded linear operators from $E_{1}$ to $E_{2}$ is denoted by $\mathcal{L}(E_{1},E_{2})$, and we abbreviate this to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$. The identity operator is indicated by $I$. Given $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ is used to refer to the real interpolation space constructed via the $K$-method, as detailed in Chapter 1 of [22]. The conjugate exponent of each $p\in[0,\infty]$ is denoted by $p^{\prime}$. The symbol $\mathcal{O}$ represents a bounded convex domain in $\mathbb{R}^{3}$ with a boundary $\partial\mathcal{O}$ that belongs to the $C^{2,0}$ class (see Section 6.2 of [11]). For $q\in[1,\infty]$, we abbreviate $L^{q}(\mathcal{O})$ to $L^{q}$, and use $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ to denote two standard Sobolev spaces (see, e.g., [34]). Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a right-continuous filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined. The expectation operator associated with this probability space is denoted by $\mathbb{E}$. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is defined by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H.$
For $p\in(1,\infty)$ and a Banach space $X$, the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$ is denoted by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$. For any Banach space $X$, the notation $\gamma(H,X)$ represents the space of all $\gamma$-radonifying operators from $H$ to $X$, as defined in [13, Chapter 9]. For a comprehensive treatment of the theory of stochastic integration with respect to the $H$-cylindrical Wiener process $W_{H}$, we refer the reader to [24].
@@
Definition of the Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, which satisfy the following conditions for every $n\in\mathbb{N}$:
$f_{n}(x,0)=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
$\lvert{f_{n}(x,y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
$\lvert{f_{n}(x,y_{1})-f_{n}(x,y_{2})}\rvert\leqslant C_{F}\lvert{y_{1}-y_{2}}\rvert,$ $\forall x\in\mathcal{O},\,\forall y_{1},y_{2}\in\mathbb{R},$ (2.3)
$\lvert{f_{n}(x_{1},y)-f_{n}(x_{2},y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert)\lvert{x_{1}-x_{2}}\rvert,$ $\forall x_{1},x_{2}\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.4)
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$, and $y_{2}$. Hereafter, with a slight abuse of notation, we will also employ $f_{n}$ to denote the Nemytskii operator defined by
$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\forall v\in L^{1}.$
Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying the summability condition
$\sum_{n\in\mathbb{N}}\lambda_{n}<\infty.$ (2.5)
For any $v\in L^{2}$, we define $F(v)\in\gamma(H,L^{2})$ as
$F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v).$ (2.6)
Accordingly, the model equation (1.1) can be explicitly formulated as follows:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0\leqslant t\leqslant T,\\ y(0)=v.\end{cases}$
Mild Solutions. Let $p,q\in[2,\infty)$. We introduce the deterministic convolution operator $S_{0}$ acting on functions $g\in L^{p}(0,T;L^{q})$ as follows:
$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad\text{for }0\leqslant t\leqslant T.$ (2.7)
Additionally, for any process $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, we define the stochastic convolution operator $S_{1}$ by
$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad\text{for }0\leqslant t\leqslant T.$ (2.8)
Here, $e^{t\Delta}$, for $t\in[0,\infty)$, denotes the analytic semigroup generated by the Laplace operator $\Delta$ subject to homogeneous Dirichlet boundary conditions. A process $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ is called a mild solution to equation (1.1) if it satisfies the equality
$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.9)
where the function $G$ is given by
$G(t):=e^{t\Delta}v,\quad\text{for }t\in[0,T].$ (2.10)
Note that the conditions in the above definition of a mild solution can be relaxed (see, e.g., [29, Chapter 6]). However, the present formulation is adequate for the purposes of this study.
Spatial Semidiscretization. Consider $\mathcal{K}_{h}$ to be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, composed of three-dimensional simplexes. We introduce the finite element space $X_{h}$ as
$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\bigm{|}v_{h}=0\text{ on }(\mathcal{O}\setminus\mathcal{O}_{h})\cup\partial\mathcal{O}_{h},\text{ and }v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\Bigr{\}},$
where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on the closure of $\mathcal{O}$, $\mathcal{O}_{h}$ is the closure of the union of the simplexes in $\mathcal{K}_{h}$, $r\geqslant 1$ is a positive integer, and $P_{r}(K)$ is the space of polynomials on simplex $K$ up to degree $r$. Let the spatial mesh size $h$ be the maximum diameter of the elements in $\mathcal{K}_{h}$. Define $P_{h}$ as the $L^{2}$-orthogonal projection operator onto $X_{h}$. The discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ is defined by
$\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\text{for all }u_{h},v_{h}\in X_{h},$
where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$. We consider the following spatial semidiscretization:
@@
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (2.11a)
$y_{h}(0)=P_{h}v,$ (2.11b)
where $v$ is the given initial value. Assume that $v\in L^{2}$. Since this spatial semidiscretization is a finite-dimensional stochastic differential equation (SDE), we can use the classical theory of finite-dimensional SDEs (see, e.g., [28, Theorem 3.21]) to conclude that Eq. 2.11 has a unique strong solution.

### v8

@@ 1 Introduction
Let $\mathcal{O}$ be a bounded, convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator equipped with homogeneous Dirichlet boundary conditions. We consider the stochastic Allen–Cahn equation given by
$\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
$y(0)=y_{0},$ (1.1b)
@@
where $y_{0}$ is the given initial condition, $W_{H}$ is an $H$-cylindrical Brownian motion, and the operator $F$ will be specified later.
@@
The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [4, 11, 18, 26, 29]. They are also an important class of nonlinear stochastic partial differential equations from a mathematical perspective. Recently, researchers in the field of numerical analysis have shown considerable interest in this topic; see [2, 3, 5, 6, 10, 17, 22, 23, 25, 26, 28, 35] and the references therein. We briefly summarize some closely related works of numerical analysis of three-dimensional stochastic Allen-Cahn equations as follows. Feng et al. [10] analyzed finite element approximations of a stochastic Allen-Cahn equation with gradient-type multiplicative noise. Majee and Prohl [25] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative Wiener process. Bréhier et al. [5] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive Wiener process. Jentzen and Pusnik [17] proposed a numerical method for approximating a class of semilinear stochastic evolution equations with non-globally Lipschitz continuous nonlinearities. Qi and Wang [28] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive Wiener process. Liu and Qiao [23] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations with multiplicative noise.
The motivation for this work stems from two important gaps in the existing literature:
  • To the best of our knowledge, the numerical analysis of the three-dimensional stochastic Allen–Cahn equation with non-smooth initial data remains underdeveloped, despite its practical significance. Existing studies often impose restrictive regularity assumptions on the initial data. For instance, the analysis in [23] requires that the initial condition $y_{0}$ belongs to the space $\dot{H}^{\alpha,2}$ for some $\alpha\in[1,2]$, while [28] assumes $\dot{H}^{1,2}$-regularity; the precise definition of these spaces is given in Section 2. The treatment of non-smooth initial data introduces substantial theoretical and numerical challenges. On the one hand, there is a lack of rigorous results for the three-dimensional stochastic Allen–Cahn equation with multiplicative noise and rough initial conditions (see, e.g., [20, 21]). On the other hand, the presence of singularities due to non-smooth initial data significantly complicates the numerical analysis.
@@
  • Current numerical studies on the three-dimensional stochastic Allen–Cahn equation with multiplicative noise predominantly focus on error estimates in the $L^{2}$-norm in space. However, the analysis of numerical approximations in more general spatial $L^{q}$-norms (for $q\geqslant 2$) remains largely unaddressed, which limits the scope and applicability of the theoretical results.
The main contributions of this work are summarized as follows:
@@
  • For the model problem Eq. 1.1 with non-smooth initial data $y_{0}\in L^{\infty}(\mathcal{O})$, we establish essential regularity estimates within the general framework of spatial $L^{q}$-norms. These estimates are crucial for the subsequent numerical analysis. Specifically, we show that the mild solution satisfies
    $y\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q}(\mathcal{O})))\quad\text{for all }p,q\in(2,\infty).$
@@
    Our analysis combines key tools from both stochastic and deterministic PDE theory, including: stochastic integration and Itô’s formula in UMD Banach spaces [8]; stochastic maximal $L^{p}$-regularity estimates [32, 33]; the Lyapunov framework for stochastic partial differential equations [20, 21]; and maximal $L^{p}$-regularity estimates for deterministic evolution equations [36].
  • To address the singularity induced by the non-smooth initial data, we employ a decomposition technique that separates the error into a singular and a more regular component. This allows us to derive the error estimate
@@
    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2}(\mathcal{O}))}\leqslant c\big{(}h^{2/p-\epsilon}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],\,p\in(2,\infty),$
    where $y_{h}$ is the strong solution of the spatial semidiscretization, $\epsilon>0$ can be arbitrarily small, and $c>0$ is a constant independent of $t$ and the spatial mesh size $h$. This result can be viewed as a stochastic counterpart of the classical deterministic error estimate given in [31, Theorem 3.2].
  • For smooth initial data $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we establish the optimal error estimate
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q}(\mathcal{O}))}\leqslant ch^{2},$
    and the nearly optimal pathwise uniform error estimate
@@
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{O}))}\leqslant ch^{2-\epsilon}\quad\forall\epsilon>0,$
    where $p,q\in(2,\infty)$ and $c>0$ is a generic constant independent of the spatial mesh size $h$. The analysis combines classical finite element theory [1, 31] with discrete maximal $L^{p}$-regularity [12] and discrete stochastic maximal $L^{p}$-regularity [19].
To the best of our knowledge, this work provides the first result on the spatial convergence of numerical approximations for the three-dimensional stochastic Allen–Cahn equation with multiplicative noise, established within the general framework of spatial $L^{q}$-norms. Although we assume deterministic initial data, the analysis can be naturally extended to the case of random initial conditions. The main contribution of this study lies in developing a rigorous and general numerical framework for spatial semidiscretizations of nonlinear stochastic parabolic equations under the framework of general spatial $L^{q}$-norms, thereby distinguishing it from previous works.
@@
The remainder of this paper is structured as follows. Section 2 introduces notation, defines the operator $F$, presents the mild solution to the model problem Eq. 1.1, and describes its spatial semidiscretization. Section 3 states the main results of this work, analyzing the regularity properties of the mild solution under non-smooth initial data, and establishes the convergence of the spatial semidiscretization for both non-smooth and smooth initial conditions. Section 4 provides the rigorous proofs of these results. Finally, Section 5 concludes the paper with a summary of the key findings.
@@ 2 Preliminaries
Notational Conventions. For Banach spaces $E_{1}$ and $E_{2}$, we denote by $\mathcal{L}(E_{1},E_{2})$ the space of all bounded linear operators from $E_{1}$ to $E_{2}$; in the case $E_{1}=E_{2}$, we write $\mathcal{L}(E_{1})$ for brevity. The identity operator on any given space is denoted by $I$. For $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ refers to the real interpolation space defined via the $K$-method, as presented in Chapter 1 of [24].
Let $\mathcal{O}\subset\mathbb{R}^{3}$ denote a bounded convex domain with $\mathcal{C}^{2}$-regular boundary $\partial\mathcal{O}$, as defined in [37, p. 39]. For $q\in[1,\infty]$, we adopt the abbreviation $L^{q}:=L^{q}(\mathcal{O})$ for Lebesgue spaces and denote by $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ the standard Sobolev spaces, following the conventions established in [14, Chapter 7]. The Laplace operator $\Delta$ is equipped with homogeneous Dirichlet boundary conditions. For $\alpha\geqslant 0$ and $q\in(1,\infty)$, define $\dot{H}^{\alpha,q}$ as the domain of the operator $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the natural norm:
$\|v\|_{\dot{H}^{\alpha,q}}:=\left\|(-\Delta)^{\alpha/2}v\right\|_{L^{q}},\quad v\in\dot{H}^{\alpha,q}.$
The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$, where $q^{\prime}$ is the Hölder conjugate of $q$. When $\alpha=2$, $\dot{H}^{2,q}$ coincides with $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$, and its norm is equivalent to the $W^{2,q}(\mathcal{O})$-norm (cf. [14, Theorem 9.15]). For $\alpha=1$, $\dot{H}^{1,q}$ agrees with $W_{0}^{1,q}(\mathcal{O})$, with equivalent norms (see [37, Theorem 16.14]).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space endowed with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$ satisfying the usual conditions. On this stochastic basis, we define a sequence $(W_{n})_{n\in\mathbb{N}}$ of independent real-valued $\mathbb{F}$-adapted Brownian motions. The associated expectation operator is denoted by $\mathbb{E}$. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is given by
$W_{H}(t)x:=\sum_{n\in\mathbb{N}}W_{n}(t)(x,h_{n})_{H},\quad\forall x\in H.$
We refer the reader to [34] for the stochastic integration with respect to $W_{H}$.
Let $p\in(1,\infty)$ and let $E$ be a Banach space. Denote by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);E)$ the closure in the Bochner space $L^{p}(\Omega\times(0,T);E)$ of the linear span of the set
[…]
where $\mathbbm{1}_{(s,t]\times F}$ denotes the indicator function of the set $(s,t]\times F$. Let $L^{0}_{\mathbb{F}}(\Omega;C([0,T];E))$ denote the space of all continuous, $\mathbb{F}$-adapted, $E$-valued processes. Two processes $g_{1}$ and $g_{2}$ in $L_{\mathbb{F}}^{0}(\Omega;C([0,T];E))$ are considered identical if $g_{1}=g_{2}$ in $C([0,T];E)$ holds $\mathbb{P}$-almost surely. We denote by $L^{p}_{\mathbb{F}}(\Omega;C([0,T];E))$ the subspace of all processes $g\in L^{0}_{\mathbb{F}}(\Omega;C([0,T];E))$ satisfying
$\mathbb{E}\big{[}\|g\|_{C([0,T];E)}^{p}\big{]}<\infty,$
where $C([0,T];E)$ is the space of continuous $E$-valued functions on $[0,T]$, equipped with the supremum norm. The space $\gamma(H,E)$ denotes the space of $\gamma$-radonifying operators from the Hilbert space $H$ to $E$ (see, e.g., [16, Chapter 9]). If $E$ is itself a Hilbert space, then $\gamma(H,E)$ is isometrically isomorphic to the space of Hilbert–Schmidt operators from $H$ to $E$.
@@
Definition of the Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of real-valued, continuously differentiable functions defined on $\overline{\mathcal{O}}\times\mathbb{R}$, satisfying the following conditions:
$\sum_{n\in\mathbb{N}}|f_{n}(x,0)|^{2}=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
$\sum_{n\in\mathbb{N}}\big{(}|f_{n}(x,y)|^{2}+|\nabla_{x}f_{n}(x,y)|^{2}\big{)}\leqslant C_{F}(1+|y|^{2}),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
$\sum_{n\in\mathbb{N}}|\nabla_{y}f_{n}(x,y)|^{2}\leqslant C_{F},$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.3)
where $C_{F}>0$ is a constant independent of $x$ and $y$. Here and in the sequel, we slightly abuse notation by identifying each $f_{n}$ with the corresponding Nemytskii operator defined pointwise by
$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\text{for all }v\in L^{2}.$
Recall that $(h_{n})_{n\in\mathbb{N}}$ denotes a fixed orthonormal basis of the separable Hilbert space $H$. Given condition Eq. 2.2, for each $v\in L^{2}$, define the operator $F(v)\in\gamma(H,L^{2})$ by
$F(v):=\sum_{n\in\mathbb{N}}h_{n}\otimes f_{n}(v),$ (2.4)
where the elementary tensor $h_{n}\otimes f_{n}(v)\in\gamma(H,L^{2})$ acts on $x\in H$ via
$(h_{n}\otimes f_{n}(v))(x):=(x,h_{n})_{H}f_{n}(v).$
In this setting, the model equation (1.1) takes the explicit form of the stochastic evolution equation:
@@
$\begin{cases}\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+\displaystyle\sum_{n\in\mathbb{N}}f_{n}(y(t))\,\mathrm{d}W_{n}(t),\quad t\in[0,T],\\ y(0)=y_{0},\end{cases}$
where $\Delta$ denotes the Laplace operator with homogeneous Dirichlet boundary conditions.
Mild Solutions. We define the two convolution operators:
$(S_{0}g)(t)$ $:=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad 0\leqslant t\leqslant T,$ (2.5)
$(S_{1}g)(t)$ $:=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad 0\leqslant t\leqslant T,$ (2.6)
whenever the integrals are well-defined. Here, $(e^{t\Delta})_{t\geqslant 0}$ denotes the analytic semigroup generated by the Laplace operator $\Delta$ under homogeneous Dirichlet boundary conditions.
Assume that $y_{0}:(\Omega,\mathcal{F}_{0},\mathbb{P})\to L^{4}$ is strongly measurable. A process $y\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$ is called a mild solution to equation (1.1) if it satisfies, $\mathbb{P}$-almost surely for all $t\in[0,T]$,
$y(t)=G(t)+\big{[}S_{0}(y-y^{3})\big{]}(t)+\big{[}S_{1}F(y)\big{]}(t),$ (2.7)
where
$G(t):=e^{t\Delta}y_{0},\quad t\in[0,T].$ (2.8)
Spatial Semidiscretization. Let $\mathcal{K}_{h}$ be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, consisting of three-dimensional simplices. We define the finite element space $X_{h}$ as
$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\Bigm{|}v_{h}=0\text{ on }\overline{\mathcal{O}}\setminus\mathcal{O}_{h},\text{ and }v_{h}\text{ is piecewise linear with respect to each }K\in\mathcal{K}_{h}\Bigr{\}},$
where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on $\overline{\mathcal{O}}$, and $\mathcal{O}_{h}$ is the closure of the union of all elements in $\mathcal{K}_{h}$. We assume that every boundary vertex of $\mathcal{O}_{h}$ lies on $\partial\mathcal{O}$. The spatial mesh size $h$ is defined as the maximum diameter over all elements in $\mathcal{K}_{h}$. Let $P_{h}$ denote the $L^{2}$-orthogonal projection onto $X_{h}$, and define the discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ via the identity
$\int_{\mathcal{O}}(\Delta_{h}u_{h})v_{h}\,\mathrm{d}x=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}x,\quad\text{for all }u_{h},v_{h}\in X_{h}.$
We consider the following spatial semidiscretization of the model equation:
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3}\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t\in[0,T],$ (2.9a)
$y_{h}(0)=P_{h}y_{0}.$ (2.9b)
Suppose that $y_{0}:(\Omega,\mathcal{F}_{0},\mathbb{P})\to L^{2}(\mathcal{O})$ is strongly measurable. A process $y_{h}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];X_{h}))$, where $X_{h}$ is equipped with the $L^{2}$-norm, is called a strong solution to (2.9) if, $\mathbb{P}$-almost surely, the identity
$y_{h}(t)=y_{h}(0)+\int_{0}^{t}\left(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3}\right)(s)\,\mathrm{d}s+\int_{0}^{t}P_{h}F(y_{h}(s))\,\mathrm{d}W_{H}(s)$
holds for all $t\in[0,T]$. By the standard stability estimate for $P_{h}$ (see, e.g., Eq. 4.19), together with the growth and Lipschitz conditions on $F$ given in Lemma 4.5, the classical theory of stochastic differential equations (see, e.g., [27, Theorem 3.21]) ensures the existence and uniqueness of a strong solution $y_{h}$ to the spatial semidiscretization (2.9).